OpenStudy (poetryprincess):

Which equation is quadratic in form? A. 2(x + 5)^2 + 8x + 5 + 6 = 0 B. X^6 + 6x^4 + 8 = 0 C. 7x^6 + 36x^3 + 5 = 0 D. 4x^9 + 20x^3 + 25 = 0

OpenStudy (omarbirjas):

try factoring out A....

OpenStudy (omarbirjas):

try factoring out A....

OpenStudy (poetryprincess):

How do I do that?

OpenStudy (omarbirjas):

put it in its simplest form. First \[2(x+5)=2x+10\], then square. \[(2x+10)=4x^2+20x+25\] now add all the like terms.\[4x^2+20x+8x+25+5+6=0\]

OpenStudy (poetryprincess):

Ok